# Qipeng Kuang Ondřej Kuželka Yuanhong Wang Yuyi Wang

### v1

@@ 4.3 Applications
Strong connectedness axiom
The strong connectedness axiom to the binary predicate $R$, denoted by $SC(R)$, requires $G(R)$ to be a strongly connected directed graph.
@@
The fragment of extended $\text{C}^{2}$ with a single strong connectedness axiom is domain-liftable. More concretely, for any extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, it holds
@@
By Lemma 2.2 we know that $\bar{\chi}_{G(\mu_{R})}(-1)=(-1)\cdot\chi_{acyc(G(\mu_{R}))}(1)$. If $G(\mu_{R})$ is strongly connected, $acyc(G(\mu_{R}))$ contains no edge, hence $\bar{\chi}_{G(\mu_{R})}(-1)=(-1)\cdot\chi_{acyc(G(\mu_{R}))}(1)=-1$. If $G(\mu_{R})$ is weakly connected but not strongly connected, $acyc(G(\mu_{R}))$ contains at least one edge so it is impossible to strictly color all vertices by one color. In this case, $\bar{\chi}_{G(\mu_{R})}(-1)=(-1)\cdot\chi_{acyc(G(\mu_{R}))}(1)=0$.
@@
A tournament is an orientation of an undirected complete graph. We can further require $G(R)$ to be strongly connected by adding the strongly connected tournament axiom to the binary predicate $R$, denoted by $SCT(R)$. For any extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, the sentence
@@
Therefore, the fragment of extended $\text{C}^{2}$ with a single strongly connected tournament axiom is domain-liftable.
Acyclicity axiom
The acyclicity axiom to the binary predicate $R$, denoted by $AC(R)$, requires that $G(R)$ does not have any directed cycle, or equivalently, $G(R)$ is a directed acyclic graph (DAG).
@@
The fragment of extended $\text{C}^{2}$ with a single acyclicity axiom is domain-liftable. More concretely, for any extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, it holds
@@
Directed tree and forest axioms
The directed tree axiom to the binary predicate $R$ and the unary predicate $Root$, denoted by $DT(R,Root)$, requires that $G(R)$ is a DAG with a single source, and the single source interprets $Root$ as true. The directed forest axiom to the binary predicate $R$, denoted by $DF(R)$, requires that $G(R)$ is a DAG such that every weakly connected component is a directed tree. In other words, $G(R)$ is a DAG such that every vertex has at most one incoming edge.
@@
The fragment of extended $\text{C}^{2}$ with a single directed tree axiom or a single directed forest axiom is domain-liftable.
@@
The proof directly follows from [MBS23, Proposition 14 and 15] that reduce WFOMC of extended $\text{C}^{2}$ with a single directed tree axiom or a single directed forest axiom to WFOMC of extended $\text{C}^{2}$ with a single acyclicity axiom. ∎
Linear order axiom
The linear order axiom to the binary predicate $R$, denoted by $LO(R)$, requires that $R$ represents a linear order. That is to say, $G(R)$ is an acyclic tournament.
@@
The fragment of extended $\text{C}^{2}$ with a single linear order axiom is domain-liftable.
@@
Theorem 4.2 together with Equation 7 yields the following equation:
$(-1)^{n}\cdot g_{n}(0,-2;\Psi_{TN},w,\overline{w},R)=\mathsf{WFOMC}(\Psi\land LO(R),n,w,\overline{w}).$
@@
Obtaining Directed Chromatic Polynomials
With the same idea of obtaining the Tutte polynomials for the graphs encoded by a fixed $\text{C}^{2}$ sentence with cardinality constraints and unary evidence, we can recover the strict and non-strict directed chromatic polynomials from SSCP and NSCP respectively of such sentences encoding directed graphs.
@@
Let $S_{1},S_{2},\cdots$ be an infinite sequence of sets of directed graphs where each graph in $S_{n}$ has $n$ vertices. Suppose that there is a fixed $\text{C}^{2}$ sentence $\Psi$ with a distinguished binary relation $E$ such that for each $n$ and each $D$ in $S_{n}$, there is an extended $\text{C}^{2}$ sentence $\Psi_{D}$ obtained by $\Psi$ conjuncted with cardinality constraints and unary evidence, and $D$ is isomorphic to $G(\mu_{E})$ for each model $\mu\in\mathcal{M}_{\Psi_{D},n}$. Then the strict and non-strict directed chromatic polynomials of any $D$ in any $S_{n}$ can be computed as follows in time polynomial in $n$:
@@
where $\mathds{1}$ is the function that always returns 1.
@@ 5 Combinations of Axioms
An important advantage of the proposed approach over the prior works is that it can easily extend to new axioms by combining existing axioms. For instance, consider the polytree axiom that requires the interpretation of a distinguished binary predicate to be a polytree, i.e., a DAG whose underlying undirected graph is a tree. Adopting the methodology of existing research [vBK23, TK23, MBS23], one would typically delve into the literature on polytree enumeration and attempt to adapt the techniques (if they exist) to the WFOMC problem. However, even if such techniques are available, integrating them into current WFOMC algorithms is not a trivial task. In contrast, our approach leverages a single polynomial to encapsulate the structural characteristics of the models. When computing WFOMC with an axiom, the process involves extracting the necessary information from the polynomials, either by algebraically manipulating the coefficients or by evaluating the polynomial at specific points. Thus, to support a new axiom that is expressible as a combination of existing axioms, we can directly and synthetically apply the existing retrieval techniques. This streamlined approach eliminates the need for further algorithmic development.
Indeed, we have demonstrated examples of axioms combination in Remark 3.1 and Remark 4.1, where the WFOMC problems present substantial challenges that existing methods have not yet addressed. The domain-liftability of these new axioms can be easily demonstrated through the fast calculation of their dependent polynomials.
@@
The fragment of extended $\text{C}^{2}$ with a single axiom in the following list is domain-liftable.
@@
  • $forest_{k}(R)=connected_{k}(R)\land forest_{k}(R)$, requiring $G(R)$ to be a forest with $k$ trees;
@@
Consider an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. We prove the theorem by showing that the WFOMC of the sentence with each of the axioms can be derived from WCP or SSCP. Let $\Psi_{R}=\Psi\land\forall x\forall y(R(x,y)\to R(y,x))\land\forall\lnot R(x,x)$,
@@
$\land\forall x\forall y\ \left(((P_{1}(x)\land P_{1}(y))\lor(P_{2}(x)\land P_{2}(y)))\to(\lnot R(x,y)\land\lnot R(y,x))\right),$
@@
6 Multi-relation Weak Connectedness Polynomial
In this section, we provide an extension of WCP, called Multi-relation Weak Connectedness Polynomial (MWCP), to further illustrate the benefits of handling axioms through the lens of polynomials. Unlike the aforementioned polynomials that are defined on a single binary relation, MWCP supports multiple binary predicates, which will allow us to handle more complex axioms on more than one predicate. Such axioms are mainly motivated by statistical relational learning (SRL), where the problems often involve multiple relations. For example, modeling the rumors spreading in a social network requires two relations: the friendship relation and the rumor-spreading relation.
The definition of MWCP shares a similar idea with WCP, using distinct variables for each relation.
Definition 6.1 (Multi-relation Weak Connectedness Polynomial).
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and unary evidence, $w$ and $\overline{w}$ be two weighting functions, and $R_{1},R_{2},\cdots,R_{k}$ be distinguished binary relations. The $n$-th MWCP $F_{n}(u_{1},\cdots,u_{k})$ of $\Psi$ for the relations $R_{1},\cdots,R_{k}$ is the unique $k$-variate polynomial which satisfies:
$F_{n}(u_{1},\cdots,u_{k};\Psi,w,\overline{w},R_{1},\cdots,R_{k})=\mathsf{WFOMC}(\Psi_{R_{1},\cdots,R_{k},u_{1},\cdots,u_{k}},n,w,\overline{w})$
for all positive integers $u_{1},\cdots,u_{k}$. Here, $\Psi_{R_{1},\cdots,R_{k},u_{1},\cdots,u_{k}}$ is defined to be
$\Psi\wedge\bigwedge_{i=1}^{k}$ $\bigwedge_{j=1}^{u_{i}}\forall x\forall y\left(A^{i}_{j}(x)\wedge(R_{i}(x,y)\lor R_{i}(y,x))\to A^{i}_{j}(y)\right)$
$\wedge$ $\bigwedge_{j=2}^{u_{i}}\forall x\left(A^{i}_{j}(x)\to A^{i}_{j-1}(x)\right),$
where $A^{i}_{j}$’s are fresh unary predicates and $w(A^{i}_{j})=\overline{w}(A^{i}_{j})=1$.
There is also a connection between MWCP and the number of connected components of the graphs $G(R_{1}),\cdots,G(R_{k})$.
Corollary 6.1 (Corollary from Proposition 3.1).
For the $n$-th MWCP of the sentence $\Psi$, the weighting functions $w$ and $\overline{w}$ and the binary relations $R_{1},\cdots,R_{k}$, it holds that
$F_{n}(u_{1},\cdots,u_{k};\Psi,w,\overline{w},R_{1},\cdots,R_{k})=\sum_{\mu\in\mathcal{M}_{\Psi,n}}W(\mu,w,\overline{w})\cdot\prod_{i=1}^{k}(u_{i}+1)^{cc(\mu_{R_{i}})},$
where $cc(\mu_{R_{i}})$ is the number of weakly connected components of $G(\mu_{R_{i}})$.
Computing MWCP might be challenging, but if the multiple relations are nested, i.e.,
$\forall x\forall y\left(R_{i}(x,y)\to R_{i+1}(x,y)\right)$
for a sequence of binary relations $R_{1},\cdots,R_{k}$, we can still compute it efficiently as shown in the following proposition.
Proposition 6.1.
Given an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and a sequence of nested binary relations $R_{1},\cdots,R_{k}$, $F_{n}(u_{1},\cdots,u_{k};\Psi,w,\overline{w},R_{1},\cdots,R_{k})$ can be computed in polynomial time in $n$.
Proof.
By Corollary 6.1, the $n$-th MWCP is a polynomial where the degree of each variable is at most $n$. The algorithm follows from the fast calculation for SCPs in Section 4.2. We index the 1-types of $\Psi_{R_{1},\cdots,R_{k},u_{1},\cdots,u_{k}}$ by $k+1$ indices $(i_{0},\cdots,i_{k})$ where $i_{0}$ represents the index of the original 1-type and others stand for the index of newly introduced predicates. Then we can observe that there are $k+1$ cases for the the 1-type coefficient $r_{(i_{0},\cdots,i_{k}),(j_{0},\cdots,j_{k})}$: for each $0\leq k^{\prime}\leq k$ such that
  • $i_{k^{\prime}}\neq j_{k^{\prime}}$,
  • for each $k^{\prime}<l\leq k$, $i_{l}=j_{l}$,
$r_{(i_{0},\cdots,i_{k}),(j_{0},\cdots,j_{k})}$ will be $\mathsf{WMC}(\psi_{i_{0},j_{0}}\land\lnot R_{1}(a,b)\land\lnot R_{1}(b,a)\land\cdots\land\lnot R_{k^{\prime}}(a,b)\land\lnot R_{k^{\prime}}(b,a))$. The new 1-type structure is a $k$-dimensional hyperspace. We can follow a similar DP strategy as Algorithm 2 but use $k$ for-loops for $u_{1},\cdots,u_{k}$, and do DP on the corresponding dimension in the order. Finally, a $k$-dimensional interpolation is used to build the polynomial from $(n+1)^{k}$ evaluations. ∎
We can show that several useful axioms in SRL applications are domain-liftable by the above proposition. The same #CC spanning subgraph axiom is an example of such axioms. The same #CC spanning subgraph axiom to the binary predicates $R_{1},R_{2}$, denoted by $SS(R_{1},R_{2})$, requires:
  • $R_{1},R_{2}$ to be symmetric;
  • $G(R_{1})$ to be a spanning subgraph of $G(R_{2})$;
  • $G(R_{2})$ to have the same number of connected components with $G(R_{1})$.
Proposition 6.2.
The fragment of extended $\text{C}^{2}$ with a single same #CC spanning subgraph axiom is domain-liftable.
Proof.
Consider an extended $\text{C}^{2}$ sentence $\Psi$, a domain of size $n$, weighting functions $w,\overline{w}$ and binary predicates $R_{1},R_{2}$. From Corollary 6.1 we have
$\mathsf{WFOMC}(\Psi\land SS(R_{1},R_{2}),n,w,\overline{w})=\sum_{k=1}^{n}[u_{1}^{k}u_{2}^{k}]F_{n}(u_{1}-1,u_{2}-1;\Psi^{\prime},w,\overline{w},R_{1},R_{2}),$
where
$\Psi^{\prime}=\Psi\land\forall x\forall y\left((R_{1}(x,y)\to R_{1}(y,x))\land(R_{2}(x,y)\to R_{2}(y,x))\right)\land\forall x(\lnot R_{1}(x,x)\land\lnot R_{2}(x,x))$
requiring $R_{1}$ and $R_{2}$ to be symmetric and irreflexive.
Since $G(R_{1})$ is a spanning subgraph of $G(R_{2})$, $R_{1}$ and $R_{2}$ are nested. By Proposition 6.1, the MWCP $F_{n}(u_{1},u_{2};\Psi,w,\overline{w},R_{1},R_{2})$ can be computed in polynomial time in $n$. ∎
With the domain-liftability of the same #CC spanning subgraph axiom, we can express some interesting constraints in SRL applications. For instance, in the rumor-spreading example, a rumor spreads through the friendship relation, and the axiom $SS(friends,spreading)$ guarantees that the rumor will spread to every connected component of the social network.
7 Experiments
We conducted experiments to verify the correctness and efficiency of our algorithms. The experiments were run on a Ubuntu server with CPU Intel Xeon Silver 4108 (1.80GHz) and 28GB RAM. The algorithms were implemented in Python. [^10] The source codes are in https://github.com/l2l7l9p/Polynomials-for-WFOMC.
7.1 Performance
We tested the runtime of computing WCP and NSCP. Since NSCP and SSCP are computed by almost the same algorithm, there is no need to test both. We used the following sentences:
  • The sentence encoding an arbitrary simple undirected graph:
    $\Psi_{ug}=\forall x\lnot R(x,x)\land\forall x\forall y\left(R(x,y)\to R(y,x)\right),$
    which is almost the simplest sentence with only one valid 1-type.
  • The sentence encoding an arbitrary simple directed graph:
    $\Psi_{dg}=\forall x\lnot R(x,x),$
    which is also a simple sentence with only one valid 1-type.
  • The classic friends-smokers sentence:
    $\Psi_{fs}=$ $\forall x\lnot fr(x,x)$
    $\land\forall x\forall y\left(fr(x,y)\to fr(y,x)\right)$
    $\land\forall x\forall y\left(fr(x,y)\land sm(x)\to sm(y)\right),$
    indicating that there is an undirected friend relationship among a group of people and if $x$ smokes, his friends will smoke. The sentence has two valid 1-types.
  • The variant of $\Psi_{fs}$, in which we require that every person has at least one friend:
    $\Psi_{fs\_1f}=\Psi_{fs}\land\forall x\exists y\ fr(x,y),$
    which has four valid 1-types due to the auxiliary Skolem predicate introduced when eliminating the existential quantifier when using the Skolemization procedure from [dBMD14].
We computed WCP and NSCP for the predicate $R$ in $\Psi_{ug}$ (for WCP only), $R$ in $\Psi_{dg}$ (for NSCP only), $fr$ in $\Psi_{fs}$ and $fr$ in $\Psi_{fs\_1f}$ over the domain size $n$ from $1$ to $25$  [^11] The running times of NSCP on $\Psi_{fs\_1f}$ with $n$ greater than $21$ are longer than $10^{5}$s, for which we omit them. and the weighting function $w,\overline{w}$ such that $w(P)=\overline{w}(P)=1$ for every predicate $P$. The performance is shown in Figure 3. In all cases, the logarithm of the running time exhibits a bend as the domain size $n$ grows, deviating from the line $2^{n}$. This behavior provides evidence that the running time for WCP and NSCP is polynomial in the domain size, aligning with the domain-liftability of Algorithm 1 and Algorithm 2.
Figure 3: The running time of computing WCP (3(a)) and NSCP (3(b)) for $\Psi_{ug}$, $\Psi_{dg}$, $\Psi_{fs}$ and $\Psi_{fs\_1f}$.
7.2 Computational Results
We evaluated the performance of Algorithm 1 and Algorithm 2 on several classes of combinatorial structures shown in Table 1 and checked the results against the Online Encyclopedia of Integer Sequences (OEIS). [^12] https://oeis.org/ In the experiments, we counted the structures described below, where we use $\Psi_{ug}$ defined above to denote the sentence encoding simple undirected graphs and $\Psi_{dg}$ defined above to denote the sentence encoding simple directed graphs.

### v2

@@ 4.3 Applications
Strong connectedness axiom. The strong connectedness axiom to the binary predicate $R$, denoted by $SC(R)$, requires $G(R)$ to be a strongly connected directed graph.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single strong connectedness axiom is domain-liftable. More concretely, for any $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, it holds
@@
By Lemma 2.16 we know that $\bar{\chi}_{G(\mu_{R})}(-1)=(-1)\cdot\chi_{acyc(G(\mu_{R}))}(1)$. If $G(\mu_{R})$ is strongly connected, $acyc(G(\mu_{R}))$ contains no edge, hence $\bar{\chi}_{G(\mu_{R})}(-1)=-1$. If $G(\mu_{R})$ is weakly connected but not strongly connected, $acyc(G(\mu_{R}))$ contains at least one edge so it is impossible to strictly color all vertices by one color. In this case, $\bar{\chi}_{G(\mu_{R})}(-1)=0$.
@@
A tournament is an orientation of an undirected complete graph. We can further require $G(R)$ to be strongly connected by adding the strongly connected tournament axiom to the binary predicate $R$, denoted by $SCT(R)$. For any $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, the sentence
@@
Therefore, the fragment of $\text{C}^{2}$ with cardinality constraints and a single strongly connected tournament axiom is domain-liftable.
Acyclicity axiom. The acyclicity axiom to the binary predicate $R$, denoted by $AC(R)$, requires that $G(R)$ does not have any directed cycle, or equivalently, $G(R)$ is a directed acyclic graph (DAG).
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single acyclicity axiom is domain-liftable. More concretely, for any $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$, it holds
@@
Directed tree and forest axioms. The directed tree axiom to the binary predicate $R$ and the unary predicate $Root$, denoted by $DT(R,Root)$, requires that $G(R)$ is a DAG with a single source, and the single source interprets $Root$ as true. The directed forest axiom to the binary predicate $R$, denoted by $DF(R)$, requires that $G(R)$ is a DAG such that every weakly connected component is a directed tree. In other words, $G(R)$ is a DAG such that every vertex has at most one incoming edge.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single directed tree axiom or a single directed forest axiom is domain-liftable.
@@
The proof directly follows from [MBS23, Proposition 14 and 15] that reduce WFOMC of $\text{C}^{2}$ with cardinality constraints and a single directed tree axiom or a single directed forest axiom to WFOMC of $\text{C}^{2}$ with cardinality constraints and a single acyclicity axiom. ∎
Linear order axiom. The linear order axiom to the binary predicate $R$, denoted by $LO(R)$, requires that $R$ represents a linear order. That is to say, $G(R)$ is an acyclic tournament with a self-loop for each vertex.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single linear order axiom is domain-liftable.
@@
Let $R^{\prime}$ be a fresh binary relation with $w(R^{\prime})=\overline{w}(R^{\prime})=1$ and define $\Psi^{\prime}_{TN}$ as Equation (7) but on relation $R^{\prime}$. Define $\Psi_{LO}$ as
$\Psi_{LO}=\Psi^{\prime}_{TN}\land(\forall x\ R(x,x))\land(\forall x\forall y(\lnot Eq(x,y)\to(R(x,y)\leftrightarrow R^{\prime}(x,y)))).$
Then Theorem 4.9 yields the following equation:
$(-1)^{n}\cdot g_{n}(0,-2;\Psi_{LO},w,\overline{w},R^{\prime})=\mathsf{WFOMC}(\Psi\land LO(R),n,w,\overline{w}).$
@@
From SSCP and NSCP to Directed Chromatic Polynomials. With the same idea of obtaining the Tutte polynomial for graphs encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence, we can recover the strict and non-strict directed chromatic polynomials from SSCP and NSCP respectively of such sentences encoding directed graphs.
@@
Let $S$ be an infinite set of directed graphs. Suppose that there is a fixed $\text{C}^{2}$ sentence $\Psi$ with a distinguished binary relation $E$ such that for each $D\in S$ of size $n$, there is a $\text{C}^{2}$ sentence $\Psi_{D}$ obtained by conjuncting $\Psi$ with cardinality constraints and ground unary literals, and $D$ is isomorphic to $G(\mu_{E})$ for each model $\mu\in\mathcal{M}_{\Psi_{D},n}$. Then the strict and non-strict directed chromatic polynomials of any $D\in S$ of size $n$ can be computed as follows in time polynomial in $n$:
@@
where $\mathds{1}$ is the weighting function that always returns 1.
@@ 5 Combinations of Axioms
An important advantage of the proposed approach over the prior works is that it can easily extend to new axioms by combining existing axioms. For instance, consider the polytree axiom that requires the interpretation of a distinguished binary predicate to be a polytree, i.e., a DAG whose underlying undirected graph is a tree. Adopting the methodology of existing research [vBK23, TK23, MBS23], one would typically delve into the literature on polytree enumeration and attempt to adapt the techniques (if they exist) to the WFOMC problem. However, even if such techniques are available, integrating them into current WFOMC algorithms is not a trivial task. In contrast, our approach leverages a single polynomial to encapsulate the structural characteristics of the models. When computing WFOMC with an axiom, the process involves extracting the necessary information from the polynomials, either by algebraically manipulating the coefficients or by evaluating the polynomial at specific points. Thus, to support a new axiom that is expressible as a combination of existing axioms, we can directly and synthetically apply the existing retrieval techniques. This streamlined approach eliminates the need for further algorithmic development for those axioms that can be captured by one of the polynomials.
Indeed, we have demonstrated examples of axiom combinations in Remarks 3.12 and 4.12, where the WFOMC problems present substantial challenges that existing methods have not yet addressed. The domain-liftability of these new axioms can be easily demonstrated through the fast calculation of their corresponding polynomials.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and a single axiom in the following list is domain-liftable.
@@
  • $forest_{k}(R)=connected_{k}(R)\land forest(R)$, requiring $G(R)$ to be a forest with $k$ trees;
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. We prove the theorem by showing that the WFOMC of the sentence with each of the axioms can be derived from WCP or SSCP. Let $\Psi_{R}=\Psi\land\forall x\forall y(R(x,y)\to R(y,x))\land\forall\lnot R(x,x)$,
@@
$\land\forall x\forall y\ \left(P_{1}(x)\land P_{1}(y)\lor P_{2}(x)\land P_{2}(y)\to\lnot R(x,y)\land\lnot R(y,x)\right),$
@@
6 Numerical Results
We implemented Algorithm 1 and Algorithm 2 and used them on several classes of combinatorial structures shown in Table 1 and checked the results against the Online Encyclopedia of Integer Sequences (OEIS). [^11] https://oeis.org/ In the evaluation, we counted the structures described below, where we use
$\Psi_{ug}=\forall x\lnot R(x,x)\land\forall x\forall y\left(R(x,y)\to R(y,x)\right)$
to denote the sentence encoding simple undirected graphs and
$\Psi_{dg}=\forall x\lnot R(x,x)$
to denote the sentence encoding simple directed graphs. We use the function $\mathds{1}$ that always returns 1 as the weighting functions.